OpenStudy (anonymous):

If the product of n and 17 is 340. What is the sum of n and 17? Thank you.

OpenStudy (ranga):

n * 17 = 340 divide by 17 n = 340 / 17 = 20 What is n + 17 if n = 20?

OpenStudy (anonymous):

37. Thanks so much